# and Reggio Emilia, Italy

### v2

@@ Abstract.
In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine. To do so, we employ 4 real-world datasets and pipelines with heterogeneous features, covering a variety of scenarios. The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand. In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best options.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 28th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc. Thus, guaranteeing high data quality is fundamental to ensure the reliability of analysis and avoid undesired additional costs (Haug et al., 2011). Oftentimes data is heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015). Due to that, data scientists dedicate considerable time and resources to perform data preparation (Fernandes et al., 2023). Typically, this fundamental process involves multiple operations called preparators (Hameed and Naumann, 2020), combined into a pipeline (Furche et al., 2016), aimed at exploring (Lissandrini et al., 2018), cleaning (Ilyas and Chu, 2019), and transforming raw data into curated datasets (Terrizzano et al., 2015; Rezig et al., 2019; Geerts et al., 2020).
@@
In particular, Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., [n. d.]). This library has become so popular in the data science ecosystem that even the popularity of Python itself has been sometimes attributed to its wide adoption (Claburn, 2017). Despite its success, Pandas notoriously presents some severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021). Firstly, it has not been designed to work with large datasets efficiently. In fact, Pandas operates in a single-threaded manner, lacks support for cluster deployment, and does not implement any memory optimization mechanism (the entire data is kept in main memory until the end of the pipeline execution). To overcome these limitations, several alternative libraries have been developed and are gaining popularity among data scientists. However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
A major reason for this confusion is the lack of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario. Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons (Schmitt, 2020; Alexander, 2023; Pinner, 2023) (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature (Petersohn et al., 2021; Rehman and Elmore, 2022; Shanbhag and Chimalakonda, 2023) either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
@@
It is common for data scientists to perform countless iterations on samples of the datasets on their laptop or PC to define a proper data preparation pipeline for the task at hand. Hence, the faster the employed library is, the less idle time they have. To support data scientists in the choice of the dataframe library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin (with both Dask and Ray engine), Polars, CuDF, Vaex, and DataTable. For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library. Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries and enables the execution in a containerized environment using Docker. This process is streamlined with the help of a convenient configuration file.
To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related literature on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021). All these datasets are collected from Kaggle [^2] https://www.kaggle.com, the most popular platform for data science challenges. This allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and have been validated by the world’s largest data science community.
We identify four essential data preparation stages in the pipelines: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). Thus, we evaluate the performance of the libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
@@ 2. Dataframes
Pandas  [^4] https://github.com/pandas-dev/pandas has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets. Pandas has established itself over time as the standard library for data manipulation and analysis. The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python. Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses makes its learning extremely straightforward.
@@
To address such limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing. The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API [^5] https://modin.readthedocs.io/en/latest or describe the parallelism between their API and the Pandas API [^6] https://datatable.readthedocs.io/en/latest/manual/comparison_with_pandas.html); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support. The selected libraries are: Pandas, PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
@@
[…]
@@
PySpark [^7] https://github.com/apache/spark is the Python API for Apache Spark, which is primarily implemented in Scala. Spark allows to work with distributed collections of Java or Python objects partitioned across a cluster, known as Relational Distributed Datasets (RDDs) (Armbrust et al., 2015). PySpark dataframes are RDDs consisting of Row objects: they can be seen as relational tables, on which relational operators can be applied. Further, they support lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015). While Spark owes its popularity to the capability of processing large-scale datasets on a cluster of machines, standalone (i.e., single-machine) mode is not only supported, but also surprisingly fast in multiple scenarios, as shown in Section 4.
@@
PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs for working with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas [^8] https://github.com/databricks/koalas), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
Preparator Pandas API SparkPD SparkSQL Modin Polars CuDF Vaex DataTable
I/O load dataframe read_csv ✓✓ ✓✓ ✓✓ ✓✓
output dataframe to_csv ✓✓ ✓✓ ✓✓ ✓✓
EDA locate missing values (nulls) isna ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
locate outliers percentile ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
search by pattern str.contains ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
sort values sort ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns list columns ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get columns types dtypes ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
get dataframe statistics describe ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
query columns query ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
DT cast columns types astype ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
delete columns drop ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
rename columns rename ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
pivot pivot_table ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
calculate column using expressions apply columnwise ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
join dataframes merge ✓✓ $\circ$ ✓✓ ✓✓ $\circ$ $\circ$
one hot encoding get_dummies ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ $\circ$
categorical encoding cat.codes ✓✓ ✓✓ ✓✓ $\circ$
group dataframe groupby ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
DC change date & time format to_datetime ✓✓ ✓✓ $\circ$ ✓✓ $\circ$ $\circ$
delete empty and invalid rows dropna ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
set content case str.lower ✓✓ ✓✓ ✓✓ ✓✓ ✓✓
normalize numeric values round ✓✓ ✓✓ ✓✓ ✓✓ ✓✓ $\circ$
deduplicate rows drop_duplicates ✓✓ ✓✓ ✓✓ $\circ$ $\circ$
fill empty cells fillna ✓✓ ✓✓ $\circ$ ✓✓ ✓✓ $\circ$
replace values occurrences replace ✓✓ ✓✓ $\circ$ ✓✓ $\circ$
edit & replace cell data apply rowise ✓✓ $\circ$ ✓✓ ✓✓ ✓✓ ✓✓
@@
Modin [^9] https://github.com/modin-project/modin is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021). Modin adopts the Pandas data format as the default storage layer and employs a set of 15 core operators to simplify Pandas functions and build its own Pandas-like API. When core operators cannot handle a function, it switches to the default to Pandas mode (affecting performance due to communication costs and Pandas single-threaded nature), reverting to a partitioned Modin dataframe after completion. It also implements opportunistic evaluation (Xin et al., 2021), enabling execution based on interactions, and facilitates incremental query construction through intermediate dataframe results. Modin is designed to dynamically switch between different partition schemes (row-based, column-based, or block-based) depending on the operation. Each partition is then processed independently by the execution engine: Dask or Ray (we consider both solutions, denoted as ModinD and ModinR, respectively).
@@
Dask  [^10] https://github.com/dask/dask is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015). It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory, and its scheduler provides flexibility to the execution. While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin $\sim$90%).
@@
Ray  [^11] https://github.com/ray-project/ray is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow [^12] https://github.com/apache/arrow (a language-independent columnar memory format, recently adopted even by Pandas  [^13] https://pandas.pydata.org/docs/dev/whatsnew/v2.0.0.html) as data format (Moritz et al., 2018). Unlike Dask, Ray does not have built-in primitives for partitioned data. While the two engines serve different use cases, their primary goal is similar: optimizing resource usage by changing how data is stored and Python code is executed.
@@
Polars [^14] https://github.com/pola-rs/polars is a Python library written in Rust and built on top of arrow2 (the Rust implementation of Arrow format). The adoption of Arrow as underlying data structure confers efficient data processing capabilities to Polars, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
@@
Polars does not use an index, but each row is indexed by its integer position in the table. Further, it has developed its own Domain Specific Language (DSL) for transforming data, whose core components are Expressions and Contexts. Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: filtering, grouping/aggregation, and selection. Polars optimizes queries through early filters and projection pushdown. Moreover, it supports both eager and lazy execution. The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can run in batches, lightening the load on memory and CPU, hence allowing to process bigger datasets (even larger than memory) in a faster manner.
@@
CuDF [^15] https://github.com/rapidsai/cudf is a component of the NVIDIA RAPIDS [^16] https://rapids.ai framework. Written in C/C++ and CUDA (hence only compatible with NVIDIA GPUs), it offers a Pandas-like API to run general-purpose data science pipelines on GPUs, leveraging their computational power for accelerating data processing.
@@
Vaex [^17] https://github.com/vaexio/vaex is a Python library [^18] Since the last commit on GitHub is dated about 10 months ago, the project might not be actively maintained at the moment. designed to handle extremely large tabular datasets (such as astronomical catalogs), producing fast statistical analysis and visualization. Supported by its vaex-core extension, written in C, it achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, allowing to explore datasets exceeding the available memory.
Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns. Only a small subset of NumPy functions benefits of lazy evaluation, storing their results as computation instructions and computing them only when needed. This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage. Vaex can process random data subsets and export data in random order for efficient resource utilization. Further, it also supports multithreading for faster computation on multi-core CPUs.
DataTable [^19] https://github.com/h2oai/datatable is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).

### v3

# Evaluation of Dataframe Libraries for Data Preparation on a Single Machine

@@ Abstract.
In this paper, we evaluate the most popular Python dataframe libraries in general data preparation use cases to assess how they perform on a single machine. To do so, we employ 4 real-world datasets with heterogeneous features, covering a variety of scenarios, and the TPC-H benchmark. The insights gained with this experimentation are useful to data scientists who need to choose which of the dataframe libraries best suits their data preparation task at hand.
In a nutshell, we found that: for small datasets, Pandas consistently proves to be the best choice with the richest API; when data fits in RAM and there is no need for complete compatibility with Pandas API, Polars is the go-to choice thanks to its in-memory execution and query optimizations; when a GPU is available, CuDF often yields the best performance, while for very large datasets that cannot fit in the GPU memory and RAM, PySpark (thanks to a multithread execution and a query optimizer) proves to be the best option.
copyright: rightsretaineddoi: https://doi.org/10.48786/edbt.2025.27isbn: 978-3-89318-098-1conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
@@ 1. Introduction
Companies and organizations significantly depend on their data to drive informed decisions, such as business strategy definition, supply chain management, etc. Thus, guaranteeing high data quality is fundamental to ensure the reliability of analysis and avoid undesired additional costs (Haug et al., 2011). Data is often heterogeneous, i.e., data sources adopt different formats and conventions (e.g., different encodings, different ways to represent dates or numeric values, etc.), and are affected by quality issues, such as missing or duplicate values (Fan, 2015). Due to that, data scientists dedicate considerable time and resources to perform data preparation (Fernandes et al., 2023). Typically, this fundamental process involves multiple operations called preparators (Hameed and Naumann, 2020), combined into a pipeline (Furche et al., 2016), aimed at exploring (Lissandrini et al., 2018), cleaning (Ilyas and Chu, 2019), and transforming raw data into curated datasets (Terrizzano et al., 2015; Rezig et al., 2019; Geerts et al., 2020).
@@
Pandas (McKinney, 2010) is by far the most widely adopted library for manipulating dataframes, and is considered by many as the de facto standard for all preparators (Petersohn et al., 2020; Jindal et al., 2021; Psallidas et al., 2022; Lee et al., 2021). This library has become so popular in the data science ecosystem that even the popularity of Python itself has been sometimes attributed to its wide adoption (Claburn, 2017).
Despite its success, Pandas notoriously presents multiple severe limits (Emani et al., 2024; Kläbe and Hagedorn, 2021). Firstly, it has not been designed to work with large datasets efficiently. In fact, Pandas operates in a single-threaded manner, lacks support for cluster deployment, and does not implement any memory optimization mechanism (the entire data is kept in main memory until the end of the pipeline execution). To overcome these limitations, several alternative libraries have been developed and are gaining popularity among data scientists. However, these libraries come with heterogeneous features (e.g., lazy evaluation, GPU support, etc.), making it hard for data scientists to navigate among them and choose the most suitable solution.
In particular, data scientists lack the support of extensive and rigorous studies to evaluate the performance of such libraries in the data preparation scenario. Public wisdom about dataframe libraries is mostly scattered across several not peer-reviewed sources doing mutual comparisons (Schmitt, 2020; Alexander, 2023; Pinner, 2023) (whose major claims are mostly confirmed by our evaluation), while related examples in scientific literature (Petersohn et al., 2021; Rehman and Elmore, 2022; Shanbhag and Chimalakonda, 2023) either focus on orthogonal aspects or only cover small subsets of libraries and preparators.
@@
It is common for data scientists to perform countless iterations on samples of the datasets on their laptop or PC to define a proper data preparation pipeline for the task at hand. Hence, the faster the employed library is, the less idle time they have. To support data scientists in the choice of the library that can speedup their workflow, we present the first extensive experimental comparison of Pandas and its most popular dataframe-based alternatives, namely PySpark, Modin, Polars, CuDF, Vaex, and DataTable. For the sake of reproducibility we developed Bento, a framework that provides a general interface to design and deploy the data preparation pipeline with any library. Bento  [^1] https://github.com/dbmodena/bento allows the definition of a common Pandas-like API for all libraries and enables the execution in a containerized environment using Docker. This process is streamlined with the help of a convenient configuration file [^2] https://github.com/dbmodena/bento/tree/master#write-a-test-file.
To cover a variety of scenarios and to provide results that are representative of real-world use cases, we consider four datasets with heterogeneous size and features, previously adopted in related work on data preparation (Hameed and Naumann, 2020) and dataframe libraries (Petersohn et al., 2021). All these datasets are collected from Kaggle (Kaggle, 2024), the most popular platform for data science challenges. This allows us to exploit published data preparation pipelines that have proven to work well for downstream machine learning tasks and have been validated by the world’s largest data science community. Moreover, we further validate our findings by testing the different libraries with the TPC-H benchmark (Pöss and Floyd, 2000).
We identify four essential data preparation stages: input/output (I/O), exploratory data analysis (EDA), data transformation (DT), and data cleaning (DC). Thus, we evaluate the performance of the libraries both on a single preparator and on sequences of preparators (i.e., the entire pipeline or its subsets corresponding to the four different stages), highlighting the impact of the query optimization techniques supported by some of them.
@@ 2. Dataframes
Pandas (Pandas, 2024a) has been the first widely-adopted Python library implementing dataframes (inspired by the corresponding data structure from the R language (McKinney, 2010)) and provides efficient and expressive data structures optimized to work with structured datasets. Pandas has established itself over time as the standard library for data manipulation and analysis. The main reasons behind its widespread popularity reside in its robust functionality and user-friendly design, making it the go-to choice for handling datasets in Python. Given its enormous diffusion, Pandas is a very mature library and the existence of a huge amount of documentation and courses makes its learning extremely straightforward.
@@
To address such limitations, many solutions have emerged to seamlessly replace Pandas in Python workflows, aiming for more efficient data processing. The selection of the Python libraries to include in our evaluation was based on three main criteria, requiring them to: (i) be based on the dataframe data structure (excluding therefore SQL-based solutions such as DuckDB (Raasveldt and Mühleisen, 2019)); (ii) be compatible with the Pandas API, to enable a fair and straightforward evaluation of each operation (note that many of these libraries, recognizing Pandas as the de facto standard, already aim at Pandas-compatibility for their API (Modin, 2024a) or describe the parallelism between their API and the Pandas API (DataTable, 2024a)); (iii) have gained more than 1k stars on GitHub, denoting the existence of a significant user base and community support. The selected libraries are: Pandas, PySpark, Modin, Polars, CuDF, Vaex, and DataTable.
@@
Native language Python Scala Python Rust C/C++ C/Python C++/Python
@@
PySpark (Apache Spark, 2024a) is the Python API for Apache Spark, which is primarily implemented in Scala.
@@
PySpark DataFrame is a type of Dataset (i.e., a distributed collection of data) organized into named columns, resembling a relational table and supporting relational operators (Apache Spark, 2024d). PySpark also supports lazy evaluation, relying on the Catalyst optimizer (Armbrust et al., 2015) and a disk spillover mechanism that automatically offloads data from RAM to disk when memory limits are reached (Apache Spark, 2024e). This allows PySpark to process datasets that exceed the machine’s physical memory capacity. While Spark owes its popularity to the capability of processing large-scale datasets on a cluster of machines, standalone (i.e., single-machine) mode is not only supported, but also surprisingly fast in multiple scenarios, as shown in Section 4.
@@
PySpark supports two different APIs: (i) Spark SQL, which allows combining SQL queries with Spark programs to work with structured data; (ii) Pandas on Spark (denoted as SparkPD and previously known as Koalas (Databricks, 2024)), which enables (by adding an index to the conventional Spark dataframe) to distribute Pandas workloads across multiple nodes without requiring modifications to the original Pandas code for most API functions ($\sim$80%).
@@
Modin (Modin, 2024b) is a Python library that provides a parallel alternative to Pandas (Petersohn et al., 2020, 2021). Modin adopts the Pandas data format as the default storage layer and employs a set of 15 core operators to simplify Pandas functions and build its own Pandas-like API. When core operators cannot handle a function, it switches to the default to Pandas mode (affecting performance due to communication costs and Pandas single-threaded nature), reverting to a partitioned Modin dataframe after completion. It also implements opportunistic evaluation (Xin et al., 2021), enabling execution based on interactions, and facilitates incremental query construction through intermediate dataframe results. Modin is designed to dynamically switch between different partition schemes (row-based, column-based, or block-based) depending on the operation. Each partition is then processed independently by the execution engine: Dask or Ray (we consider both solutions, denoted as ModinD and ModinR, respectively).
@@
Dask (Dask, 2024b) is a Python library for distributed computing, which allows working with large distributed Pandas dataframes (Rocklin, 2015). It is designed to efficiently extend memory capacity using disk space, making it suitable for systems with limited memory, and its scheduler provides flexibility to the execution. While we also considered the inclusion of Dask as an independent library, we do not report its results since we found it to perform better in combination with Modin (as Dask is not well suitable for a single machine) and it only covers $\sim$55of the Pandas API (while Modin covers $\sim$90%) (Modin, 2024c).
@@
Ray (Ray, 2024b) is a general-purpose framework for parallelizing Python code, using an in-memory distributed storage system and Apache Arrow (Apache Arrow, 2024a) (a language-independent columnar memory format, recently adopted even by Pandas (Pandas, 2023)) as data format (Moritz et al., 2018). Unlike Dask, Ray does not have built-in primitives for partitioned data. While the two engines serve different use cases, their primary goal is similar: optimizing resource usage by changing how data is stored and Python code is executed.
@@
Polars (Polars, 2024d) is a Python library written in Rust and built on top of arrow2, the Rust implementation of the Arrow format. The adoption of Arrow as the underlying data structure provides Polars with efficient data processing capabilities, optimized through parallel execution, cache-efficient algorithms, and efficient usage of resources.
@@
Polars does not use an index, but each row is indexed by its integer position in the table. Further, it has developed its own Domain Specific Language (DSL) for transforming data, whose core components are Expressions and Contexts. Expressions facilitate concise and efficient data transformations, while Contexts categorize evaluations into three main types: filtering, grouping/aggregation, and selection. Polars optimizes queries through early filters and projection pushdown. Moreover, it supports both eager and lazy evaluation. The lazy strategy allows to run queries in a streaming manner: instead of processing the entire data at once, they can run in batches, lightening the load on memory and CPU, hence allowing to process bigger datasets (even larger than memory) in a faster manner.
@@
CuDF (NVIDIA, 2024a) is a component of the NVIDIA RAPIDS framework. Written in C/C++ and CUDA (hence only compatible with NVIDIA GPUs), it offers a Pandas-like API to run general-purpose data science pipelines on GPUs, leveraging their computational power for accelerating data processing.
@@
Vaex (Vaex, 2024c) is a Python library [^4] Since the last commit on GitHub is dated about one year ago, the project might not be actively maintained at the moment. designed to handle extremely large tabular datasets (such as astronomical catalogs), producing fast statistical analysis and visualization. Supported by its vaex-core extension, written in C, Vaex achieves memory optimization through streaming algorithms, memory-mapped files, and a zero-copy policy, supporting the exploration of datasets that exceed the available memory.
Vaex enables efficient column-wise operations by wrapping a series of NumPy arrays as columns. Only a small subset of NumPy functions benefits of lazy evaluation, storing their results as computation instructions and computing them only when needed. This is achieved through virtual columns, expressions that can be added to datasets without incurring additional memory usage. Vaex can process random data subsets and export data in random order for efficient resource utilization. Moreover, it also supports multithreading for faster computation on multi-core CPUs.
DataTable (H2O, 2024) is a Python package that relies on the Frame object, similar to Pandas dataframes or SQL tables (i.e., data arranged in a two-dimensional array with rows and columns).